a student releases a marble from the top of the ramp, the marble increases speed steadily and travels 90.0cm at 4.90s. what is the marbles final speed?
1
Expert's answer
2020-10-12T07:47:00-0400
"s=0.5at^2\\a=\frac{v}{t}\\s=0.5vt\\0.9=0.5(4.9)v\\v=0.367\frac{m}{s}"